OpenStudy (anonymous):

Solve the inequality- Negative x over six is less than or equal to 3.

OpenStudy (anonymous):

\[\frac{ -x }{6 }\le3\]

OpenStudy (anonymous):

Can you solve that? that's how it was originally written, but i didn't know how to put it in fraction form

OpenStudy (anonymous):

multiply both sides by x -x <= (3)*(6) -x<=18 // now the tricky part is that when you divide by -1 over an inequality, you have to reverse the < to > so we have \[x \ge -18\]

OpenStudy (anonymous):

Thanks :D